A 91-kg parent and a 21-kg child meet at the centre of an ice rink. They place their hands together and push. (a) Is the force experienced by the child more than, less than, or the same as the force experienced by the parent? (b) Is the acceleration ...
Is the acceleration of the child more than, less than, or the same as the acceleration of the parent?
1
Expert's answer
2020-03-19T11:30:30-0400

(a) The third Newton's law states


"\bf F_{12}=-\bf F_{21}."

Hence, the force experienced by the child is the same as the force experienced by the parent.

(b) The second Newton's law states


"a_1=\frac{F_{21}}{m_1}, \: a_2=\frac{F_{12}}{m_2}."

or


"\frac{a_{\rm child}}{a_{\rm parent}}=\frac{m_{\rm parent}}{m_{\rm child}}=\frac{91}{21}"

Hence, the acceleration of the child is more than the acceleration of the parent.